please i need help with this function f(x)=x^2+5, find f(a+h)-f(a)
Answer by stanbon(75887)

f(x)=x^2+5, find f(a+h)-f(a)
--------
Caluculate each term separately:
f(a+h)=(a+h)^2+5
=a^2+2ah+h^2+5
f(a)=a^2+5
--------
Subtract f(a) from f(a+h) to get:
2ah+h^2
=========
Cheers,
Stan H.